# Rethinking cluster-conditioned diffusion models

### v1

@@ Abstract
We present a comprehensive experimental study on image-level conditioning for diffusion models using cluster assignments. We elucidate how individual components regarding image clustering impact image synthesis across three datasets. By combining recent advancements from image clustering and diffusion models, we show that, given the optimal cluster granularity with respect to image synthesis (visual groups), cluster-conditioning can achieve state-of-the-art FID (i.e. 1.67, 2.17 on CIFAR10 and CIFAR100 respectively), while attaining a strong training sample efficiency. Finally, we propose a novel method to derive an upper cluster bound that reduces the search space of the visual groups using solely feature-based clustering. Unlike existing approaches, we find no significant connection between clustering and cluster-conditional image generation. The code and cluster assignments will be released.
Keywords:
cluster conditioning conditional image generation conditional image synthesis
@@ 1 Introduction
Diffusion models have recently enabled enormous progress in many visual generative tasks such as image synthesis [35, 44, 66] and manipulation [91]. Conditioning diffusion models with human-annotated data is today’s standard practice as it significantly improves the image fidelity [29, 11, 7]. Image-level conditioning is typically achieved using annotated pairs from text captions [75, 77] or class labels [35, 62]. Empirical and theoretical evidence shows that label-conditional diffusion models are faster to train and outperform unconditional ones by leveraging large amounts of image-label pairs [7]. Human-annotated labels often contain inaccuracies [13, 9], while publicly available image-text pairs can be non-descriptive [12].
@@
Figure 1: An ideal image-level conditioning shall group together images solely based on visual cues, shown in the same row, which do not always align with human labels, indicated above each image (CIFAR100[51] samples).
Large, human-annotated datasets are prohibitively costly [69] and hard to obtain in a plethora of real-life applications such as medical image synthesis [47]. Human annotators classify images (group visual stimuli into discrete concepts) based on non-visual cues [68], which may originate, for example, from language learning [53, 10]. As a result, a dataset can have different annotation levels and hierarchies; a wolf can have the label wolf, animal, mammal, or carnivore. To investigate the image misalignment with the human-derived labels (ground truth) and its impact on conditional generative learning, we focus on cluster-based conditioning on diffusion models. Unsupervised image classification or image clustering refers to algorithmically assigning a semantic group to an image [58], called a cluster, given an a priori number of such groups [17, 21, 43]. This way, the clustering accuracy can be computed using the linear one-to-one cluster-to-label mapping estimated from the Hungarian method [52]. Such a limitation is not required for generative learning, and an optimal cluster alignment with ground truth (GT) labels does not always guarantee optimal cluster-conditional image generation (Fig. 1).
Clusters are computed solely based on visual input, which renders them a promising candidate for alternative image-level conditioning. Apart from allowing us to access the quality of human labels for conditional image synthesis, clusters enable us to leverage the benefits of conditioning on unsupervised datasets without human annotations. An ideal image-level conditioning signal would be based solely on shared characteristics corresponding to general visual concepts, as illustrated in Fig. 1. To this end, a clustering algorithm would sometimes group images into more fine-grained groups than the GT labels, i.e.,  sub-grouping analog and digital devices that map to the same label, such as telephone, or merge similarly looking images with different labels (merging different types of trees).
Deep image clustering has recently seen significant progress [89, 92, 3], especially using off-the-shelf features learned from large-scale pre-training [87, 2, 27]. In particular, the learned feature representations from visual self-supervision capture informative features by learning to match augmented views of the same image [23, 31, 18] or by masking parts of the input [8, 32].
@@
It is currently unclear how clustering performance relates to cluster-conditional image synthesis performance for multiple reasons. First, clustering metrics, such as clustering accuracy, measure the alignment to GT labels, while generative models might benefit more from visually aligned groups [39]. Second, clustering metrics cannot be computed for unlabeled datasets, their primary use case. Third, existing metrics cannot fairly evaluate cluster assignments with different numbers of clusters. They cannot, therefore, help determine the optimal cluster granularity w.r.t. generative performance for an arbitrary dataset, which has not been sufficiently explored [7, 57].
@@
Existing cluster-conditioned generative approaches [57, 7, 39] use simple generative baselines [35] and have mainly investigated $k$-means clustering on balanced classification datasets. The adoption of $k$-means is justified due to its ability to work with off-the-shelf image features and non-parametric nature, apart from setting the number of clusters. Conversely, it is well-established that $k$-means is suboptimal for image clustering because its cluster assignments are highly imbalanced [87, 39]. Still, the image generation performance’s strong sensitivity to the number of clusters [7] and the choice of pre-trained model [39] has been established. Even though recent feature-based image clustering methods have demonstrated superior performance compared to $k$-means, their applicability to conditional image generation has not been investigated.
In this paper, we systematically study clustering as an alternative condition to ground truth labels for diffusion models using recent advancements from image clustering and diffusion methods [2, 44]. Our key contributions are summarized as follows:
  • We show that, given the optimal cluster granularity, cluster-conditioning can achieve state-of-the-art FID (i.e. 1.67, 2.17 on CIFAR10 and CIFAR100, respectively) while attaining a strong training sample efficiency.
@@
  • We propose a novel and computationally efficient method to derive an upper cluster bound using solely feature-based clustering. As a result, we reduce the search space of the cluster granularity for optimal conditional image synthesis performance. We validate the discovered upper bound on three datasets and two generative metrics.
  • Finally, we find no significant correlation between clustering performance and the associated cluster-conditional generative performance. This lack of correlation is observed across clustering methods, performance metrics, and pre-trained models for feature extraction.
@@ 2 Related Work
2.1 Conditional generative models using human labels.
@@
Early attempts on controllable image synthesis adopted generative adversarial networks (GANs) [61]. By conditioning both the generator and discriminator on human labels, GANs can produce images of a specific GT label, even at the scale of ImageNet [16]. Recently, diffusion models (DMs) have emerged as an expressive and flexible category of generative models [79, 80, 36], delivering unprecedented sample quality even at high resolutions [73, 46, 12]. Internal guidance-based methods [30, 36, 37, 38] further improved the flexibility of DMs during sampling such as classifier and classifier-free guidance, which can significantly improve visual fidelity, similar to the truncation trick for GANs [16]. Another family of methods focuses on reducing the sampling steps of DMs, namely distillation-based DMs [76, 81, 94, 59, 56]. DMs exhibit an enormous number of label-conditional variants [35, 49, 44, 62, 67]. Recently, Stein et al. [83] proposed the adoption of Fréchet DINOv2 [63] distance as it aligns better with human preferences and demonstrated that DMs achieve the highest perceptual realism, based on human evaluations.
2.2 Alternative conditioning of generative models.
@@
Unlike internal guidance [37, 36, 38], external conditioning signals are computed from the training set using additional models, which is commonly referred to as self-conditioning. These models are typically pre-trained with self-supervision [18] or natural language supervision [69]. Self-conditioning signals can be roughly divided into image-level and sub-image-level. Image-level conditioning refers to a single condition for all pixels in an image, such as cluster assignments or text captions [73]. Sub-image-level conditions refer to specific parts or regions of an image [39, 6]. For instance, Hu et al. [39] extend DMs to incorporate conditioning of bounding boxes or segmentation masks.
Feature-based conditioning. Using image features or features from the images’ nearest neighbor (NN) set can also provide an informative conditioning signal. Bordes et al. [15] reveal how different self-supervised representations impact visual fidelity in DMs. Instance-conditioned GAN combines GT labels with features from each image’s NN set [19]. Sheynin et al. [78] train a diffusion model on unlabeled data using the nearest neighbor captions from a labeled image-text reference dataset.
@@
CLIP-based conditioning. The well-aligned multi-modal semantic space of contrastive language image pre-trained models (CLIP [69]) enabled many new types of conditioning. Ramesh et al. [71] leverage the zero-shot capability of CLIP to condition DMs on language-guided image representations to improve diversity. Zhou et al. provide a language-free framework [97] that directly generates text features for images using CLIP. Still, image caption models are necessary to derive a descriptive text caption for conditional DMs [12]. In contrast, CLIP models require an a priori set of descriptive candidate captions, such as label names.
Cluster-based conditioning. Similar to label conditioning, conditioning on cluster assignments facilitates DMs by allowing them to specialize on a distinct set of visual features shared between image groups [39, 7]. To address the limited diversity of GANs on unlabeled datasets, the adoption of online $k$-means clusters was proposed for GAN training [57]. In subsequent work, Bao et al. [7] computed the $k$-means clusters offline using contrastive learned features on the training data. However, their approach could not outperform the label-conditioned models.
@@
The first cluster-based approach that achieves competitive performance compared to GT labels leverages pre-trained feature extractors [39], which is the closest to our work. In [39], the authors attempt to establish a correlation between clustering and cluster-conditional generative performance for different self-supervised backbones. The observed correlation was fairly weak, cannot be tested on unlabeled datasets, and its sensitivity to the number of clusters is unknown. While it was demonstrated that DINO [18] provides the most informative $k$-means clusters [39], their method lacks a strategy for choosing the cluster granularity for unlabeled datasets. Finally, the authors used the denoising diffusion probabilistic model (DDPM [35]) instead of more recent state-of-the-art DMs, which is likely the reason why they do not obtain competitive performance.
@@ 3.1 Notations and prerequisites
We consider the frequent scenario where we have access to an unlabeled dataset $D$ and a pre-trained self-supervised feature extractor $g(\cdot)$. We denote the number of ground truth (GT) labels as $C_{GT}$, which we assume to be unknown. To distinguish cluster-based conditioning from the image clustering task, we denote the number of visual groups $C_{V}$ of $D$ as the optimal number of clusters w.r.t. image synthesis (e.g. as measured by FID [34]). We adopt the diffusion model approach of Karras et al. (EDM [44]) as a baseline, which introduced various improvements to the standard DDPM [35] in terms of network architecture, training objective, sampling, and data augmentation strategies [45]. We provide an overview of EDM in the supplementary. Although we focus on EDM throughout this work, our method can be applied to any conditional generative model.
@@
TEMI clustering. Given an a priori determined number of clusters $C$ and a feature extractor $g(.)$, TEMI [2] first mines the $m$ nearest neighbors (NN) of all ${\mathbf{x}}\in D$ in the feature space of $g(\cdot)$ based on their cosine similarity. We denote the set of NN for ${\mathbf{x}}$ by $S_{{\mathbf{x}}}$. During training, TEMI randomly samples ${\mathbf{x}}$ from $D$ and ${\mathbf{x}}^{\prime}$ from $S_{{\mathbf{x}}}$ to generate image pairs with similar (visual) features. A self-distillation framework is introduced to learn the cluster assignments with a teacher and student head $h_{t}(\cdot)$ and $h_{s}(\cdot)$ that share the same architecture (i.e. 3-layer MLP) but differ w.r.t. their parameters. The features of the image pair are fed to the student and teacher heads $h^{i}_{s}({\mathbf{z}})$, $h^{i}_{t}({\mathbf{z}}^{\prime})$, where $i\in\{1,\dots,H\}$ is the head index and ${\mathbf{z}}=g({\mathbf{x}})$, ${\mathbf{z}}^{\prime}=g({\mathbf{x}}^{\prime})$. The outputs of the heads are converted to probabilities $q_{s}(c|{\mathbf{x}})$ and $q_{t}(c|{\mathbf{x}}^{\prime})$ using a softmax function. The TEMI objective maximizes the pointwise mutual information between images $x$ and $x^{\prime}$, using the clustering index $c$ as information bottleneck [43]
@@
$\mathcal{L}_{\text{\sc{temi}}}^{i}({\mathbf{x}},{\mathbf{x}}^{\prime}):=-\frac{1}{H}\sum_{j=1}^{H}\sum_{c^{\prime}=1}^{C}q_{t}^{j}(c^{\prime}|{\mathbf{x}})q_{t}^{j}(c^{\prime}|{\mathbf{x}}^{\prime})\log\sum_{c=1}^{C}\frac{\left(q_{s}^{i}(c|{\mathbf{x}})q_{t}^{i}(c|{\mathbf{x}}^{\prime})\right)^{\gamma}}{\tilde{q}_{t}^{i}(c)},$ (1)
@@
Here, $\tilde{q}_{t}^{i}(c)$ is an estimate of the teacher cluster distribution $\mathbb{E}_{{\mathbf{x}}\sim p_{\text{data}}}\left[q_{t}^{i}(c|{\mathbf{x}})\right]$, which can be computed by an exponential moving average over mini-batches $B$ defined as $\tilde{q}_{t}^{i}(c)\leftarrow\lambda\,\tilde{q}_{t}(c)+(1-\lambda)\frac{1}{\absolutevalue{B}}\sum_{{\mathbf{x}}\in B}q_{t}^{i}(c|{\mathbf{x}})$.
@@
The hyperparameter $\gamma\in(0.5,1]$ avoids the collapse of all sample pairs into a single cluster while $\lambda\in(0,1)$ is a momentum hypeparameter. The loss function Eq.(1) is further symmetrized and averaged over heads to obtain the final training loss. After training, the cluster assignments $c^{*}({\mathbf{x}})=\operatorname*{arg\,max}_{c}{q_{t}(c|{\mathbf{x}})}$ and the empirical cluster distribution $q(c)$ on the training set are computed from the TEMI head with the lowest loss. We use $c^{*}({\mathbf{x}})$ as a condition of the generative model during training and sample from $q(c)$ to generate images. Adaloglou et al. [2] show that the value $\gamma=0.6$ enforces a close to uniform cluster utilization and achieves state-of-the-art clustering accuracy for $C_{GT}$[2].
3.2 Cluster-conditional EDM (C-EDM) diffusion model

### v2

@@ Abstract
Diffusion-based image generation models can enhance image quality when conditioned on ground truth labels. Here, we conduct a comprehensive experimental study on image-level conditioning for diffusion models using cluster assignments. We investigate how individual clustering determinants, such as the number of clusters and the clustering method, impact image synthesis across three different datasets. Given the optimal number of clusters with respect to image synthesis, we show that cluster-conditioning can achieve state-of-the-art performance, with an FID of 1.67 for CIFAR10 and 2.17 for CIFAR100, along with a strong increase in training sample efficiency. We further propose a novel empirical method to estimate an upper bound for the optimal number of clusters. Unlike existing approaches, we find no significant association between clustering performance and the corresponding cluster-conditional FID scores. Code is available at https://github.com/HHU-MMBS/cedm-official-wavc2025
@@ 1 Introduction
Diffusion models have enabled significant progress in many visual generative tasks, such as image synthesis [29, 39, 57] and manipulation [80]. Conditioning diffusion models on human-annotated data is today’s standard practice as it significantly improves the image fidelity [23, 8, 25]. Image-level conditioning is typically realized by using associated text captions [65, 66] or class labels, if available [29, 53]. Nonetheless, human-annotated labels are costly and often contain inaccuracies [10, 7], while publicly available image-text pairs can be non-descriptive [9].
@@
Figure 1: An ideal image-level conditioning should group images based on shared patterns, shown in the same row, which do not always align with human labels, indicated above each image (CIFAR100[45] samples).
Large, human-annotated datasets are prohibitively costly [59] and hard to obtain in a plethora of real-life applications such as medical image synthesis [41]. Human annotation is subject to the label collection procedure and can have varying annotation hierarchies [31, 11, 3]. An ideal conditioning signal would be based solely on shared characteristics corresponding to general visual concepts, as illustrated in Fig. 1. To provide a ground truth (GT) label-free alternative that is applicable to unlabelled datasets, we focus on cluster-based conditioning on diffusion models. Image clustering refers to algorithmically assigning a semantic group to an image, called a cluster, given an a priori number of such groups [14, 18, 38].
Maximizing cluster alignment with ground truth (GT) labels does not guarantee optimal cluster-conditional image generation. To create visually distinguishable groups (Fig. 1), a clustering algorithm would sometimes group images into more fine-grained groups than the GT labels, i.e.,  sub-grouping analog and digital devices that map to the same label or merge similarly looking images with different labels (merging different types of trees).
@@
Deep image clustering has recently seen significant progress [78, 81, 4], especially when utilizing off-the-shelf features learned from large-scale pre-training [76, 2, 3]. Currently, the relationship between clustering performance and cluster-conditional image synthesis performance remains unclear for several reasons. First, clustering metrics capture the alignment to GT labels, whereas generative models may benefit more from visually distinguishable groups [34]. Second, clustering metrics cannot be computed for unlabeled datasets. Third, existing metrics are not suitable for fairly evaluating cluster assignments with varying numbers of clusters. Therefore, they cannot help to determine the optimal cluster granularity w.r.t. generative performance for an arbitrary dataset. This aspect has not been sufficiently explored to date [50, 25].
@@
Existing cluster-conditioned generative approaches [50, 25, 34] often adopt simple generative baselines [29] and have mainly investigated $k$-means clustering on balanced classification datasets. The adoption of $k$-means is justified due to its ability to work with off-the-shelf feature extractors and its non-parametric nature, apart from setting the number of clusters. Conversely, it is well-established that $k$-means is suboptimal for image clustering because its cluster assignments are highly imbalanced [76, 34]. While cluster-conditioned approaches are sensitive to the number of clusters [25], there is currently no method to predetermine this number. Previous methods have overlooked this [50, 25, 34], significantly hindering their applicability to unlabeled datasets. Recent feature-based image clustering methods have demonstrated superior performance compared to $k$-means [76, 2], and their efficacy in conditional image generation is yet to be investigated. For the above reasons, cluster-conditioning methods underperform GT conditional ones to date.
@@
In this paper, we systematically study clustering as a label-free condition for diffusion models. We demonstrate that optimal cluster granularity achieves state-of-the-art FID scores (1.67 for CIFAR10 and 2.17 for CIFAR100) while enhancing training sample efficiency. We propose a computationally efficient method to derive an upper cluster bound using feature-based clustering, which narrows the search space for optimal image synthesis performance. We validate this upper bound across three datasets and two generative metrics. Finally, we find no significant association between clustering performance and cluster-conditional image synthesis performance across various clustering methods, performance metrics, and pre-trained models.
@@ 2 Related Work
2.1 Conditional generative models
@@
Early attempts on controllable image synthesis adopted generative adversarial networks (GANs) [52]. By conditioning both the generator and discriminator on human labels, GANs can produce images for a specific GT label, even at the scale of ImageNet [13]. Recently, diffusion models (DMs) have emerged as an expressive and flexible category of generative models [68, 69, 30, 63, 40]. Internal guidance-based methods [24, 30, 32, 33] further improved the flexibility and visual fidelity of DMs during sampling such as classifier and classifier-free guidance. DMs exhibit an enormous number of label-conditional variants [29, 44, 39, 53, 58]. Recently, Stein et al. [71] proposed the adoption of Fréchet DINOv2 [54] distance as it aligns better with human preferences and demonstrated that DMs achieve the highest perceptual realism.
2.2 Alternative conditioning of generative models
@@
Unlike internal guidance [32, 30, 33], external conditioning signals are computed from the training set using additional models, which is commonly referred to as self-conditioning. Self-conditioning signals can be roughly divided into image-level and sub-image-level. Image-level conditioning refers to a single condition for all pixels in an image, such as cluster assignments or text captions [63]. Sub-image-level conditions refer to specific parts or regions of an image [34, 6]. For instance, Hu et al. [34] extend DMs to incorporate conditioning of bounding boxes or segmentation masks.
@@
Feature-based conditioning. Using image or text features can also provide an informative conditioning signal [12, 67, 49, 61]. Bordes et al. [12] condition DMs directly on image representations (oracle features from real images) and reveal their impact on visual fidelity. Instance-conditioned GAN combines GT labels with features from each image’s NN set [16]. Ramesh et al. [61] leverage the zero-shot capability of CLIP to condition DMs on language-guided image representations to improve diversity. Zhou et al. provide a language-free framework [84] that directly generates text features for images using CLIP. Still, CLIP models require an a priori set of descriptive candidate captions, such as label names.
@@
Cluster-based conditioning. Similar to label conditioning, conditioning on cluster assignments facilitates DMs by allowing them to specialize on a distinct set of shared visual features [34, 25, 50]. In [25], Bao et al. computed the $k$-means clusters offline using contrastive learned features on the training data. However, their approach could not outperform the label-conditioned models. The first cluster-based approach that achieves competitive performance compared to GT labels leverages pre-trained feature extractors [34], which is the closest to our work. In [34], the authors attempt to weak establish a correlation between clustering and cluster-conditional generative performance for different self-supervised models. The observed correlation cannot be tested on unlabeled datasets, and its sensitivity to the number of clusters is unknown. While it was demonstrated that DINO [15] provides the most informative $k$-means clusters [34], their method lacks a strategy for choosing the cluster granularity for unlabeled datasets.
@@ 3.1 Notations and prerequisites
We consider the frequent scenario where we have access to an unlabeled dataset $D$ and a pre-trained feature extractor $g(\cdot)$. We denote the number of ground truth (GT) labels as $C_{GT}$, which we assume to be unknown. To distinguish cluster-based conditioning from the image clustering task, we denote the number of visual groups $C_{V}$ of $D$ as the optimal number of clusters w.r.t. image synthesis (e.g. as measured by FID [28]). We adopt the diffusion model approach of Karras et al. (EDM [39]) as a baseline, which introduced various improvements to the standard DDPM [29]. We provide an overview of EDM in the supplementary material. Although we focus on EDM throughout this work, our method can be applied to any conditional generative model.
@@
TEMI clustering. Given an a priori determined number of clusters $C$ and a feature extractor $g(.)$, TEMI [2] first mines the $m$ nearest neighbors (NN) of all ${\mathbf{x}}\in D$ in the feature space of $g(\cdot)$ based on their cosine similarity. We denote the set of NN for ${\mathbf{x}}$ by $S_{{\mathbf{x}}}$. During training, TEMI randomly samples ${\mathbf{x}}$ from $D$ and ${\mathbf{x}}^{\prime}$ from $S_{{\mathbf{x}}}$ to generate image pairs with similar (visual) features. A self-distillation framework is introduced to learn the cluster assignments with a teacher and student head $h_{t}(\cdot)$ and $h_{s}(\cdot)$ that share the same architecture (i.e. 3-layer MLP) but differ w.r.t. their parameters. The features of the image pair are fed to the student and teacher heads $h^{i}_{s}({\mathbf{z}})$, $h^{i}_{t}({\mathbf{z}}^{\prime})$, where $i\in\{1,\dots,H\}$ is the head index and ${\mathbf{z}}=g({\mathbf{x}})$, ${\mathbf{z}}^{\prime}=g({\mathbf{x}}^{\prime})$. The outputs of the heads are converted to probabilities $q_{s}(c|{\mathbf{x}})$ and $q_{t}(c|{\mathbf{x}}^{\prime})$ using a softmax function. The TEMI objective,
@@
$\begin{split}\mathcal{L}_{\text{\sc{temi}}}^{i}({\mathbf{x}},{\mathbf{x}}^{\prime}):=-\frac{1}{H}\sum_{j=1}^{H}\sum_{c^{\prime}=1}^{C}q_{t}^{j}(c^{\prime}|{\mathbf{x}})q_{t}^{j}(c^{\prime}|{\mathbf{x}}^{\prime})\\ \log\sum_{c=1}^{C}\frac{\left(q_{s}^{i}(c|{\mathbf{x}})q_{t}^{i}(c|{\mathbf{x}}^{\prime})\right)^{\gamma}}{\tilde{q}_{t}^{i}(c)}\quad,\end{split}$ (1)
@@
maximizes the pointwise mutual information between images ${\mathbf{x}}$ and ${\mathbf{x}}^{\prime}$, using the clustering index $c$ as information bottleneck [38]. Here, $\tilde{q}_{t}^{i}(c)$ is an estimate of the teacher cluster distribution $\mathbb{E}_{{\mathbf{x}}\sim p_{\text{data}}}\left[q_{t}^{i}(c|{\mathbf{x}})\right]$, which can be computed by an exponential moving average over mini-batches $B$ defined as $\tilde{q}_{t}^{i}(c)\leftarrow\lambda\,\tilde{q}_{t}(c)+(1-\lambda)\frac{1}{\absolutevalue{B}}\sum_{{\mathbf{x}}\in B}q_{t}^{i}(c|{\mathbf{x}})$.
@@
The hyperparameter $\gamma\in(0.5,1]$ avoids the collapse of all sample pairs into a single cluster, and $\lambda\in(0,1)$ is a momentum hyperparameter. The loss function Eq. 1 is further symmetrized and averaged over heads to obtain the final training loss. After training, the cluster assignments $c^{*}({\mathbf{x}})=\operatorname*{arg\,max}_{c}{q_{t}(c|{\mathbf{x}})}$ and the empirical cluster distribution $q(c)$ on the training set is computed from the TEMI head with the lowest loss. We use $c^{*}({\mathbf{x}})$ as a condition of the generative model during training and sample from $q(c)$ to generate images. Adaloglou et al. [2] show that the value $\gamma=0.6$ enforces a close to uniform cluster utilization and achieves state-of-the-art clustering accuracy for $C_{GT}$[2].
3.2 Cluster-conditional EDM (C-EDM)